# Reweighted Spectral Partitioning Works: Bounds for Special Graph Classes

### v3

@@ 3.1 Geometric Intersection Graphs and Nearest-Neighbour Graphs
[…]
$\sum_{i=1}^{n}V_{d}\cdot r_{i}^{\prime\prime d}\leq\sum_{i=1}^{n}\text{volume}(B_{i}^{\prime})\leq k\cdot A_{d}.$
[…]
$\dot{\gamma}^{(d)}(G)\leq\frac{\sum_{v\in V}r_{i}^{\prime\prime 2}}{n}\leq\left(\frac{\sum_{v\in V}r_{i}^{\prime\prime d}}{n}\right)^{\frac{2}{d}}\leq\left(\frac{A_{d}\cdot k}{V_{d}\cdot n}\right)^{\frac{2}{d}},$
where the second inequality follows from the power-mean inequality. ∎
Figure 2: An example of a geodesic ball $B_{i}^{\prime}$ on $S^{1}$ embedded in $\mathbb{R}^{2}$, with $c_{i}^{\prime},r_{i}^{\prime},r_{i}^{\prime\prime}$ all labelled.
As previously mentioned, we will also apply this result to $k$-nearest neighbour graphs, again in a similar manner to Spielman and Teng [ST96, ST07, Corollary 5.2]. Let $\tau_{d}$ denote the kissing number in $d$ dimensions, which is the maximum number of non-overlapping unit balls in $\mathbb{R}^{d}$ arranged to all touch a central unit ball. It is known that as $d\to\infty$, $2^{0.2075d(1+\operatorname{o}(1))}\leq\tau_{d}\leq 2^{0.401d(1+\operatorname{o}(1))}$ [Wyn65, KL78]. Hence, $\log\tau_{d}\gtrsim d$ and $\tau_{d}^{\frac{1}{d}}\lesssim 1$. Miller, Teng, Thurston, and Vavasis [MTTV97] observed that every $k$-nearest neighbour graph is the subgraph of a $\tau_{d}k$-ply neighbourhood system, and moreover that every $k$-nearest neighbour graph has maximum degree bounded by $\tau_{d}k$. Hence, we obtain the following corollary:
Corollary 3.7.
@@
$\frac{\lambda_{2}(G)}{\Delta}\leq\lambda_{2}^{*}(G)\leq\gamma^{(d+1)}(G)<2\cdot\left(2\pi\cdot\left(d+2\right)\right)^{\frac{1}{d}}\cdot\left(\frac{\tau_{d}k}{n}\right)^{\frac{2}{d}}\lesssim\left(\frac{k}{n}\right)^{\frac{2}{d}}.$
Consequently, $\psi(G)\lesssim\sqrt{\min\{d,\log\Delta\}}\cdot\left(\frac{k}{n}\right)^{\frac{1}{d}}$. Moreover, reweighted spectral partitioning will find, in polynomial time, a sparse vertex cut $S$ with $\frac{|\partial S|}{|S|}\lesssim\sqrt{\log\Delta}\cdot\left(\frac{k}{n}\right)^{\frac{1}{d}}\lesssim\sqrt{d+\log k}\cdot\left(\frac{k}{n}\right)^{\frac{1}{d}}$.
Note that the bound on $\frac{|\partial S|}{|S|}$ is simplified due to the bound $\log\Delta\lesssim\log\tau_{d}k\lesssim d+\log k$.
The above two statements apply algorithmically to reweighted spectral partitioning. We note that, in some applications, the geometric information may actually be known, in which case the constructions above can be used more directly:
Corollary 3.8.
Let $G$ be a $d$-dimensional $k$-ply neighbourhood system with maximum degree $\Delta$, provided as a set of $d$-dimensional coordinates and radii for each vertex. Then, in polynomial time, we can compute a balanced 2/32/32 / 3-vertex-separator of size $\operatorname{\mathcal{O}}\left(\sqrt{\min\{d,\log\Delta\}}\cdot\left(\frac{k}{n}\right)^{\frac{1}{d}}\right)$.
Corollary 3.9.
Let $G$ be a $d$-dimensional $k$-nearest neighbour graph with maximum degree $\Delta$, provided as a set of $d$-dimensional coordinates and a value $k$. Then, in polynomial time, we can compute a balanced 2/32/32 / 3-vertex-separator of size $\operatorname{\mathcal{O}}\left(\sqrt{\min\{d,\log\Delta\}}\cdot\left(\frac{k}{n}\right)^{\frac{1}{d}}\right)$.
@@ 3.2 Bounded-Genus Graphs
Kelner’s bound on $\lambda_{2}$ (stated in Theorem 2.3) for triangulated $g$ graphs also makes use of circle packings [Kel04, Kel06b, Kel06a], and, as we will see, can also be generalized to a bound on $\lambda_{2}^{*}$ and $\gamma^{(1)}$. However, the resulting bound also has a (large) polynomial dependence on the maximum degree $\Delta$. Rather than directly applying this bound, we will apply a new transformation that allows us to almost fully tighten this dependence. Specifically, our new transformation will take a $g$ graph with $n$ vertices and maximum degree $\Delta$, and produce a graph with $O(\Delta\cdot n)$ vertices and constant maximum degree. We will then relate the values of $\gamma^{(1)}$ for the two graphs. In the case of $\lambda_{2}^{*}$ and $\gamma^{(1)}$, this relationship will result in a greatly improved $O(\log\Delta)$ factor for the upper bound.
Recall that a graph $G$ has genus $g$ if it has a combinatorial embedding (represented as a rotation system) onto a surface of genus $g$. That is, each vertex has a circular ordering on its vertices, forming a locally-planar embedding. Such an embedding is cellular if each face induced by the rotation system is simple. A cellular embedding is a triangulation if each face defined by the rotation system is a triangle.
It is notable that Kelner’s bound on $\lambda_{2}$ does not apply to non-triangulated $g$ graphs, despite the paper claiming otherwise. The claimed argument is as follows:
  • Adding edges to a graph can only increase $\lambda_{2}$.
@@
  • Apply the bound on triangulated $g$ graphs that Kelner proves in the rest of his paper. Note that this bound relies on a generalized circle packing theorem of He and Schramm [HS93] that applies only to simple triangulations of $g$ surfaces.
@@
  • A maximal $g$ graph is not always a triangulation. In particular, $K_{5}$ is a maximal genus-111 graph that is not triangulated in any embedding into the torus. We may prove this by contradiction: Suppose this was not the case, and that it had a triangulated embedding in the torus. Then, by the Euler characteristic, such an embedding must have 555 faces. We may count the number of edge-face incidences. Since the embedding is triangulated (and hence cellular), each edge is incident to exactly two faces, and there are 101010 edges and hence 210=202$\cdot$ 10=202 ⋅ 10 = 20 incidences in total. Since the embedding is triangulated, there are exactly 333 edges incident to each face, and hence there are 35=153$\cdot$ 5=153 ⋅ 5 = 15 incidences in total – a contradiction.
  • Naively adding edges may increase the maximum degree, possibly up to $\Omega(n)$, and the bound that Kelner obtains for triangulations has a dependence on the maximum degree.
It is worth noting that neither of these issues arise for planar graphs: Maximal planar graphs are always planar triangulations, and there is a known result that triangulates a planar graph while asymptotically preserving the maximum degree [KB92, KB97]. For genus $g>0$, we will present a method to bypass these issues with carefully chosen Steiner vertices while simultaneously tightening and generalizing the bound to $\lambda_{2}^{*}$ and $\gamma^{(1)}$.
Henceforth, we will fix a rotation system for our initial graph $G$. This rotation system need not be cellular, but we will assume that $G$ is simple (as we do throughout this paper). Importantly for our purposes, this will allow us to make certain local transformations to a graph without changing its genus, so long as the rotation system is respected in these transformations. These transformations will make use of the following structure:
@@
Let $G=(V,E)$ be a graph with $n$ vertices, and let $H=(V^{\prime},E^{\prime})$ be a graph with $r\cdot n$ vertices. Suppose there is a function $p:V^{\prime}\to V$ so that $|p^{-1}(v)|=r$ for each $v\in V$, so that $p^{-1}(v)|$ is also a connected subset of vertices with diameter $L$. Suppose furthermore that for each edge $uv\in E^{\prime}$, $p(u)p(v)\in E$. Then we say that $G$ is a uniform shallow minor of $H$ with depth $L$.
Usually when studying minors, we usually care about “forbidden” minors of a particular given graph. However, we will use uniform shallow minors in the opposite sense: Given a graph $G$, we will usually aim to find a graph $H$ for which $G$ is a uniform shallow minor. In particular, their use is characterized by the following technical lemma:
@@
Let $G=(V,E)$ be a graph with $n$ vertices, and let $H=(V^{\prime},E^{\prime})$ be a graph with $|V^{\prime}|=r\cdot n$ vertices. Suppose $G$ is a uniform shallow minor of $H$ with depth $L$. Then $\gamma^{(1)}(G)\lesssim r\cdot L\cdot\gamma^{(1)}(G)$.
@@
We start with a trivial case whose elimination we will use later: If $\gamma^{(1)}(H)\cdot r\cdot L\gtrsim 1$, then the result holds since $\gamma^{(1)}(G)\lesssim 1$ (see 1.3).
[…]
Let $\pi(v)$ be a random variable that is uniformly chosen over $p^{-1}(v)$, so that each $\pi(v)$ is independent. Let $m(v)$ be samples of all such $\pi(v)$ so that
@@
Let $f_{G}(v):=f_{H}(m(v))-\frac{1}{n}\sum_{x\in V}f_{H}(m(x))$, and let $y_{G}(v):=4r\cdot(2L+1)\cdot\sum_{v^{\prime}\in p^{-1}(v)}y_{H}(v)$.
Under these choices, the objective claim is satisfied since $\sum_{v\in V}y_{G}(v)=4r\cdot(2L+1)\cdot\sum_{v^{\prime}\in V^{\prime}}y_{H}(v)$, and hence we need only check that the constraints of $\gamma^{(1)}(G)$ are satisfied by $f_{G},y_{G}$.
The first constraint is satisfied since $\sum_{v\in V}f_{G}(v)=\sum_{v\in V}f_{H}(m(v))-\sum_{x\in V}f_{H}(m(x))=\overline{0}$. Next, note that for any vertex $v\in V$, and any two vertices $u^{\prime},v^{\prime}\in p^{-1}(v)$, there is a path $u^{\prime}=v_{1}^{\prime},\dots,v_{l+1}^{\prime}=v^{\prime}$ of length $l\leq L$ contained entirely within $p^{-1}(v)$. Hence, by the triangle inequality and Cauchy-Schwarz,
$\begin{array}[]{r c l}|f_{H}(u^{\prime})-f_{H}(v^{\prime})|^{2}&\leq&\left[|f_{H}(v_{1}^{\prime})-f_{H}(v_{2}^{\prime})|+\cdots+|f_{H}(v_{l}^{\prime})-f_{H}(v_{l+1}^{\prime})|\right]^{2}\\[4.0pt] &\leq&l\left[|f_{H}(v_{1}^{\prime})-f_{H}(v_{2}^{\prime})|^{2}+\cdots+|f_{H}(v_{l}^{\prime})-f_{H}(v_{l+1}^{\prime})|^{2}\right]\\[4.0pt] &\leq&L\left[|f_{H}(v_{1}^{\prime})-f_{H}(v_{2}^{\prime})|^{2}+\cdots+|f_{H}(v_{l}^{\prime})-f_{H}(v_{l+1}^{\prime})|^{2}\right]\\[4.0pt] &\leq&L\left[y(v_{1}^{\prime})+2y(v_{2}^{\prime})+2y(v_{3}^{\prime})+\cdots+2y(v_{l}^{\prime})+y(v_{l+1}^{\prime})\right],\end{array}$
and so for a fixed $v\in V$,
$\begin{array}[]{r l}&\sum_{u^{\prime}\in p^{-1}(v)}\sum_{v^{\prime}\in p^{-1}(v)}|f_{H}(u^{\prime})-f_{H}(v^{\prime})|^{2}\\[4.0pt] \leq&\sum_{u^{\prime}\in p^{-1}(v)}\sum_{v^{\prime}\in p^{-1}(v)}L\left[y(v_{1}^{\prime})+2y(v_{2}^{\prime})+2y(v_{3}^{\prime})+\cdots+2y(v_{l}^{\prime})+y(v_{l+1}^{\prime})\right]\\[4.0pt] \leq&2L\sum_{u^{\prime}\in p^{-1}(v)}\sum_{v^{\prime}\in p^{-1}(v)}\sum_{w^{\prime}\in p^{-1}(v)}y(w^{\prime}).=2Lr^{2}\sum_{w^{\prime}\in p^{-1}(v)}y(w^{\prime})\\[4.0pt] \end{array}$
Using this bound, we can also get a lower bound on $\sum_{v\in V}|f_{G}(v)|^{2}$:
$\begin{array}[]{r c l}2|V|\cdot\sum_{v\in V}|f_{G}(v)|^{2}&=&\sum_{u\in V}\sum_{v\in V}|f_{G}(u)-f_{G}(v)|^{2}\\[12.0pt] &=&\sum_{u\in V}\sum_{v\in V}|f_{H}(m(u))-f_{H}(m(v))|^{2}\\[12.0pt] &\geq&\operatorname{\mathbb{E}}\left[\sum_{u\in V}\sum_{v\in V}|f_{H}(\pi(u))-f_{H}(\pi(v))|^{2}\right]\\[12.0pt] &=&\frac{1}{r^{2}}\sum_{u\in V}\sum_{v\in V\setminus\{u\}}\sum_{u^{\prime}\in p^{-1}(u)}\sum_{v^{\prime}\in p^{-1}(v)}|f_{H}(u^{\prime})-f_{H}(v^{\prime})|^{2}\\[12.0pt] &=&\frac{1}{r^{2}}\left[\sum_{u^{\prime}\in V^{\prime}}\sum_{v^{\prime}\in V^{\prime}}|f_{H}(u^{\prime})-f_{H}(v^{\prime})|^{2}\right.\\ &&\left.-\sum_{v\in V}\sum_{u^{\prime}\in p^{-1}(v)}\sum_{v^{\prime}\in p^{-1}(v)}|f_{H}(u^{\prime})-f_{H}(v^{\prime})|^{2}\right]\\[16.0pt] &\geq&\frac{1}{r^{2}}\left[2|V|\cdot r\cdot\sum_{v^{\prime}\in V^{\prime}}|f_{H}(v^{\prime})|^{2}\right.\\ &&\left.-2Lr^{2}\cdot\left(\sum_{v^{\prime}\in V^{\prime}}y_{H}(v^{\prime})\right)\cdot\left(\sum_{v^{\prime}\in V^{\prime}}|f_{H}(v^{\prime})|^{2}\right)\right]\\[12.0pt] &=&\frac{2\sum_{v^{\prime}\in V^{\prime}}|f_{H}(v^{\prime})|^{2}}{r}\left[|V|-r\cdot(L/2)\cdot\gamma^{(1)}(H)\right]\\[12.0pt] &\geq&\frac{2\sum_{v^{\prime}\in V^{\prime}}|f_{H}(v^{\prime})|^{2}}{r}\left[|V|-1\right],\end{array}$
where the last step follows from our early assumption on the value of $\gamma^{(1)}(H)$. Thus, $\sum_{v\in V}|f_{G}(v)|^{2}\geq\frac{\sum_{v^{\prime}\in V^{\prime}}|f_{H}(v^{\prime})|^{2}}{2r}$.
@@
It remains to show the constraint for each edge is satisfied. Let $e=uv\in E$. Similarly to our earlier discussion, there is a path of length $k\leq 2L+1$ in $H$ from $m(u)$ to $m(v)$ using only vertices in $p^{-1}(\{u,v\})$. Denote this path as $m(u)=v_{1},\dots,v_{k+1}=m(v)$. Then,
$\begin{array}[]{r c l}y_{G}(u)+y_{G}(v)&=&4r(2L+1)\sum_{u^{\prime}\in p^{-1}(u)}y_{H}(u^{\prime})+\sum_{v^{\prime}\in p^{-1}(v)}y_{H}(v^{\prime})\\[8.0pt] &\geq&4r(2L+1)\left[y_{H}(v_{1})+\cdots+y_{H}(v_{k+1})\right]\\[8.0pt] &\geq&2r(2L+1)\sum_{i=1}^{k}\left[y_{H}(v_{i})+y_{H}(v_{i+1})\right]\\[8.0pt] &\geq&2r(2L+1)\frac{\sum_{i=1}^{k}\left|f_{H}(v_{i})-f_{H}(v_{i+1})\right|^{2}}{\sum_{x\in V^{\prime}}|f_{H}(x)|^{2}}\\[8.0pt] &\geq&(2L+1)\frac{\sum_{i=1}^{k}\left|f_{H}(v_{i})-f_{H}(v_{i+1})\right|^{2}}{\sum_{x\in V}|f_{G}(x)|^{2}}\\[8.0pt] &\geq&(2L+1)\frac{\left[\sum_{i=1}^{k}\left|f_{H}(v_{i})-f_{H}(v_{i+1})\right|\right]^{2}}{k\sum_{x\in V}|f_{G}(x)|^{2}}\\[8.0pt] &\geq&(2L+1)\frac{\left|f_{H}(v_{1})-f_{H}(v_{k+1})\right|^{2}}{k\sum_{x\in V}|f_{G}(x)|^{2}}\\[8.0pt] &\geq&\frac{\left|f_{G}(u)-f_{G}(v))\right|^{2}}{\sum_{x\in V}|f_{G}(x)|^{2}},\end{array}$

### v4

@@ 5.2 Ball-Intersection Graphs and Nearest-Neighbour Graphs
Define $h,r,\phi$ as in the $2$-dimensional cross-section of the $4$-dimensional unit ball given in Figure 4.
Figure 4: Cross-section of a unit ball, and the different associated quantities.
The $3$-dimensional area of the $3$-dimensional sphere-cap with colatitude angle $\phi$ (equivalently, height $h$) is $\int_{0}^{\phi}A_{2}(\sin\theta)^{2}d\theta=\frac{A_{2}}{2}(\phi-\cos\phi\sin\phi)$. Note that $A_{2}=4\pi$ is the surface area of the unit $2$-sphere, and that $\cos\phi=1-h$, so $\phi=\arccos(1-h)$. For the volume, $r=\sqrt{2h}$, so $V_{3}\cdot r^{3}=V_{3}\cdot(2h)^{\frac{3}{2}}$. Using, say, $h=0.4$, we can compare the values, and verify that the $3$-dimensional surface area of the sphere cap is <2.81<2.81, while $V_{3}\cdot r^{3}>2.99$. ∎
A weaker form of their bound still holds, fortunately. We will prove and use the following inequality:
Lemma 5.7.
For a sphere cap $C$ on the $d$-dimensional unit sphere embedded in $\mathbb{R}^{d+1}$, let $f$ be its centre and let $r$ be the maximum Euclidean distance in $\mathbb{R}^{d+1}$ from $f$ to a point in $C$. Assume $C$ contains at most half the surface area of the sphere. Then, $d$-dimensional surface area of the sphere cap has $\text{area}(C)\geq\frac{A_{d}r^{d}}{4^{d}}$.
We defer the proof of this inequality to Appendix D, since it is a fairly straightforward expansion of some Gamma and Beta functions. Using it, we can now prove the main result of this subsection:
Proposition 5.8 (Restatement of Proposition 2.5).
Let $G$ be the intersection graph of $d$-dimensional balls with ply $k$ and maximum degree $\Delta$. Then,
$\gamma^{(d+1)}(G)\lesssim\left(\frac{k}{n}\right)^{\frac{2}{d}}.$
Proof.
The proof is very similar to that of Proposition 2.4, and again is based on the methods of Spielman and Teng [ST96, ST07, proof of Theorem 5.1]. By Theorem 5.4, we may realize $G$ as the geometric intersection graph of a set of geodesic balls on $S^{d}$, so that every point in $S^{d}$ is contained in at most $k$ of these geodesic balls almost surely, and so that the centroid of the centres of the geodesic balls in $\mathbb{R}^{d+1}$ is exactly the origin. Denote the geodesic balls (sphere caps) as $C_{1},\dots,C_{n}$. For the geodesic ball $C_{i}$, let $f_{i}$ denote its centre and let $r_{i}$ denote the maximum Euclidean distance in $\mathbb{R}^{d+1}$ from $f_{i}$ to a point in $C_{i}$. By Lemma 5.7, $\text{area}(C_{i})\geq\frac{A_{d}r_{i}^{d}}{4^{d}}$. The total sum of all such areas is bounded above by $kA_{d}$, so $\sum_{i=1}^{n}r_{i}^{d}\leq k4^{d}$. By the power-mean inequality, we get $\frac{\sum_{i=1}^{n}r_{i}^{2}}{n}\leq\left(\frac{\sum_{i=1}^{n}r_{i}^{d}}{n}\right)^{\frac{2}{d}}\leq\left(\frac{k4^{d}}{n}\right)^{\frac{2}{d}}\leq 16\left(\frac{k}{n}\right)^{\frac{2}{d}}$. ∎
As previously mentioned, we will also apply this result to $k$-nearest neighbour graphs, again in a similar manner to Spielman and Teng [ST96, ST07, Corollary 5.2]. Let $\tau_{d}$ denote the kissing number in $d$ dimensions, which is the maximum number of non-overlapping unit balls in $\mathbb{R}^{d}$ arranged to all touch a central unit ball. It is known that as $d\to\infty$, $2^{0.2075d(1+\operatorname{o}(1))}\leq\tau_{d}\leq 2^{0.401d(1+\operatorname{o}(1))}$ [Wyn65, KL78]. Hence, $\log\tau_{d}\gtrsim d$ and $\tau_{d}^{\frac{1}{d}}\lesssim 1$. Miller, Teng, Thurston, and Vavasis [MTTV97] observed that every $k$-nearest neighbour graph is the subgraph of an intersection graph of balls with ply $\tau_{d}k$, and moreover that every $k$-nearest neighbour graph has maximum degree bounded by $\tau_{d}k$. Hence, we obtain the following corollary:
Corollary 5.9 (Restatement of Corollary 2.6).
@@
$\gamma^{(d+1)}(G)\lesssim\left(\frac{k}{n}\right)^{\frac{2}{d}}.$
Note that the weaker bound on $\frac{|\partial S|}{|S|}$ stated in Corollary 1.8 for this class comes from the inequality $\log\Delta\lesssim\log\tau_{d}k\lesssim d+\log k$.
We will now give specialized algorithms for these graph classes, since the implied separator results are new.
Corollary 5.10 (Restatement of Corollary 2.10).
Let $G$ be the intersection graph of $d$-dimensional balls with ply $k$ and maximum degree $\Delta$, provided as a set of $d$-dimensional coordinates and radii for each vertex. Then, in polynomial time, we can compute a balanced $2/3$-vertex-separator of size
$\operatorname{\mathcal{O}}\left(\sqrt{\min\{d,\log\Delta\}}\cdot\left(\frac{k}{n}\right)^{\frac{1}{d}}\right).$
Proof.
Instead of applying Theorem 1.4, perform the stereographic projection implicit in the proof of Theorem 5.4, and then directly apply Theorem 4.2 with Algorithm 1, as well as Lemma 5.2 (taking the best of the two). Finally, apply Theorem 4.1 with Algorithm 4. ∎
Corollary 5.11 (Restatement of Corollary 2.11).
Let $G$ be a $d$-dimensional $k$-nearest neighbour graph with maximum degree $\Delta$, provided as a set of $d$-dimensional coordinates and a value $k$. Then, in polynomial time, we can compute a balanced $2/3$-vertex-separator of size
$\operatorname{\mathcal{O}}\left(\sqrt{\min\{d,\log\Delta\}}\cdot\left(\frac{k}{n}\right)^{\frac{1}{d}}\right).$
Proof.
We start by computing the graph itself. By using a simple construction of Kwok, Lau, and Tung [KLT22, Proof of Lemma 3.12], we can use the coordinates to obtain radii for a representation of the graph as a ball-intersection graph, while optimizing the objective function up to constant factors (relative to the fixed choice of coordinates). Then, instead of applying Theorem 1.4, perform the stereographic projection implicit in the proof of Theorem 5.4, and directly apply Theorem 4.2 with Algorithm 1, as well as Lemma 5.2 (taking the best of the two). Finally, apply Theorem 4.1 with Algorithm 4. ∎
Another result of Spielman and Teng for a graph class they call $\alpha$-overlap graphs also relied on the incorrect inequality for the area of sphere caps [ST07, Theorem 6.4]. Lemma 5.7 may also be applied to fix that result in essentially the same manner. The result could be generalized to a bound on $\gamma^{(1)}$ as well.
@@ 5.3 Bounded-Genus Graphs
In this subsection, we will prove Proposition 2.7, which states that a genus-$g$ graph $G$ with maximum degree $\Delta$ has $\gamma^{(1)}(G)\lesssim\frac{g\log\Delta}{n}$.
This will involve several steps. At a high-level, we will use two different groups of techniques:
  • Techniques involving uniform shallow-minors, which we will define shortly.
  • Techniques based on the theory of circle packings for genus-$g$ graphs, adapting constructions used by Kelner to give a bound on the Fiedler value $\lambda_{2}$ for triangulated genus-$g$ graphs [Kel04, Kel06b, Kel06a].
We will now briefly outline some important definitions and challenges related to genus-$g$ graphs. Recall that a genus-$g$ graph is one that can be drawn on a genus-$g$ (oriented) surface without crossing edges. Such a drawing is called a surface embedding, and it implies the existence of a rotation system: A (circular) ordering of the edges around each vertex. Like a plane graph, we can define a face of a genus-$g$ graph as any cycle that is formed by a tour induced by always moving counter-clockwise in the rotation system at a vertex to choose the next edge, starting from the previous. A rotation system (and a corresponding embedding) is called cellular if each edge is a part of exactly two such tours. Not every surface embedding is cellular, even for plane graphs: A face on a plane graph can include an edge twice, each time traversed in opposite directions. For instance, this is always true of cut edges in any embedding of a planar graph (see Figure 5). In a surface embedding for a genus-$g$ graph with $g\geq 1$, it can also be the case that an edge is included twice in the same direction (see Figure 5).
Figure 5: Non-cellular embeddings: planar graph (left) and $K_{5}$ in the flat torus (right). A non-cellular face in each is traced.
For simple graphs, we call a graph embedded in a surface with cellular faces a triangulation if all of its faces consist of exactly $3$ edges. Planar triangulations correspond exactly to maximal planar graphs, which are graphs for which the addition of any edge would enforce the invalidation of planarity, regardless of embedding. For graphs of genus-$g$, this is no longer true:
Example 5.12.
$K_{5}$ is a maximal graph of genus $1$, and $K_{5}$ admits no triangulated embedding into a surface of genus $1$.
See Figure 5 for an embedding of $K_{5}$ in the torus.
Proof.
The proof is by contradiction. Suppose this was not the case, and that it had a triangulated embedding in the torus. $K_{5}$ has $5$ vertices and $10$ edges. By the Euler characteristic, an embedding of $K_{5}$ in the torus must have $5$ faces. We will now count the number of edge-face incidences in two different ways. Since the embedding is triangulated (and hence cellular), each edge is incident to exactly two faces, and there are $10$ edges and hence $2\cdot 10=20$ incidences in total. Since the embedding is triangulated, there are exactly $3$ edges incident to each face, and hence there are $3\cdot 5=15$ incidences in total, giving a contradiction. ∎
Kelner claimed to have bounded $\lambda_{2}$ for all genus-$g$ graphs, rather than just triangulations. The method he used to try and prove this was to reduce the case of general genus-$g$ graphs to the case of triangulated genus-$g$ graphs. However, the reduction used for this is actually incorrect. The claimed argument is as follows:
  • Adding edges to a graph can only increase $\lambda_{2}$ (this is also true for $\gamma^{(1)}$).
@@
  • Apply the bound on triangulated genus-$g$ graphs that Kelner proves in the rest of his paper. This bound relies on a generalized circle packing theorem of He and Schramm [HS93] that applies only to simple triangulations of genus-$g$ surfaces.
@@
  • As mentioned, a maximal genus-$g$ graph is not always a triangulation.
  • Even in a planar graph, naively adding edges may increase the maximum degree, possibly up to as large as $\Omega(n)$, and the bound that Kelner obtains for triangulations has a dependence on the maximum degree.
It is worth noting that both of these issues can be sidestepped for planar graphs: Maximal planar graphs are always planar triangulations, and there is a known result that triangulates a planar graph while asymptotically preserving the maximum degree [KB92, KB97]. It is possible that these issues could be amended for genus-$g$ graphs by allowing augmentation to obtain a multigraph, but such a construction is not known, and it would need to be checked that Kelner’s construction can be made to work with a multigraph. Instead, we have both generalized and improved Kelner’s bound via our other techniques.
We now move onto the proof. This will involve several phases. Henceforth, we will fix a rotation system for our initial graph $G$. This rotation system need not be cellular, but we will assume that $G$ is simple (as we do throughout this paper). Importantly for our purposes, so long as we carefully respect this rotation system, we will be able to make certain local transformations to a graph without changing its genus.
5.3.1 Uniform Shallow Minors
For the first phase of our transformations, we will make use of the following structure:
@@ 5.3.1 Uniform Shallow Minors
Let $G=(V,E)$ be a graph with $n$ vertices, and let $H=(V^{\prime},E^{\prime})$ be a graph with $r\cdot n$ vertices. Suppose there is a function $p:V^{\prime}\to V$ so that $|p^{-1}(v)|=r$ for each $v\in V$, and $p^{-1}(v)$ is a connected subset of vertices with diameter $L$. Suppose furthermore that for each edge $uv\in E^{\prime}$, $p(u)p(v)\in E$. Then we say that $G$ is a uniform shallow minor of $H$ with depth $L$.
See Figure 6 for some examples.
Figure 6: Two examples of uniform shallow minors: A graph that admits $C_{4}$ as a uniform shallow minor with depth 22 (left), and a graph that admits $P_{4}$ as a uniform shallow minor with depth 11 (right).
Usually when studying minors, we start with a graph, and then consider which minors it contains or does not contain. Often, we even “forbid” minors, resulting in classes like $K_{h}$-minor-free graphs. However, we will use uniform shallow minors in the exact opposite sense: Given a graph $G$, we will aim to find a graph $H$ for which $G$ is a uniform shallow minor. In particular, their use is characterized by the following technical lemma:
@@
Let $G=(V,E)$ be a graph with $n$ vertices, and let $H=(V^{\prime},E^{\prime})$ be a graph with $|V^{\prime}|=r\cdot n$ vertices. Suppose $G$ is a uniform shallow minor of $H$ with depth $L$. Then $\gamma^{(1)}(G)\lesssim r\cdot L\cdot\gamma^{(1)}(H)$.
@@
Figure 7: Examples of the construction in Lemma 5.14 applied to the uniform shallow minors in Figure 6.
@@
We start with a trivial case whose elimination we will use later: If $\gamma^{(1)}(H)\cdot r\cdot L\geq 1$, then the result holds since $\gamma^{(1)}(G)\lesssim 1$: We can obtain a solution validating this bound by picking two vertices $u,v$ of the graph, and setting $y(u)=y(v)=2$, f(u)=1,f(v)=1f(u)=1,f(v)=-1, and $y(w)=f(w)=0$ for any $w\not\in\{u,v\}$.
At a high-level, our construction will randomly sample a representative vertex in $H$ for each vertex in $G$, and then look at the edges along the shortest-paths between these representatives. This is visualized in Figure 7.
Let $p:V^{\prime}\to V$ be the function that certifies the depth-$L$ uniform shallow minor. Let $f_{H},y_{H}$ be a solution to $\gamma^{(1)}(H)$. We will construct a solution $f_{G},y_{G}$ to $\gamma^{(1)}(G)$ with random sampling as follows: Let $\pi(v)$ be a random variable that is uniformly chosen over $p^{-1}(v)$, so that each $\pi(v)$ is independent. Let $m(v)$ be samples of all such $\pi(v)$ so that
@@
Let $f_{G}(v):=f_{H}(m(v))-\frac{1}{|V|}\sum_{x\in V}f_{H}(m(x))$, and let $y_{G}(v):=2(2L+1)\cdot\sum_{v^{\prime}\in p^{-1}(v)}y_{H}(v^{\prime})$.
@@
We check that the constraints of $\gamma^{(1)}(G)$ are satisfied by $f_{G},y_{G}$. The normalization constraint is satisfied since $\sum_{v\in V}f_{G}(v)=\sum_{v\in V}f_{H}(m(v))-\sum_{x\in V}f_{H}(m(x))=\overline{0}$. For the edge constraints, let $e=uv\in E$. There is a path of length $k\leq 2L+1$ in $H$ from $m(u)$ to $m(v)$ using only vertices in $p^{-1}(\{u,v\})$. Denote this path as $m(u)=v_{1},\dots,v_{k+1}=m(v)$. Then,
$\begin{array}[]{r c l}y_{G}(u)+y_{G}(v)&=&2(2L+1)\left[\sum_{u^{\prime}\in p^{-1}(u)}y_{H}(u^{\prime})+\sum_{v^{\prime}\in p^{-1}(v)}y_{H}(v^{\prime})\right]\\[8.0pt] &\geq&2(2L+1)\left[y_{H}(v_{1})+\cdots+y_{H}(v_{k+1})\right]\\[8.0pt] &\geq&(2L+1)\sum_{i=1}^{k}\left[y_{H}(v_{i})+y_{H}(v_{i+1})\right]\\[8.0pt] &\geq&(2L+1)\sum_{i=1}^{k}\left|f_{H}(v_{i})-f_{H}(v_{i+1})\right|^{2}\\[8.0pt] &\geq&\frac{2L+1}{k}\left[\sum_{i=1}^{k}\left|f_{H}(v_{i})-f_{H}(v_{i+1})\right|\right]^{2}\\[8.0pt] &\geq&\frac{2L+1}{k}\left|f_{H}(v_{1})-f_{H}(v_{k+1})\right|^{2}\\[8.0pt] &=&\frac{2L+1}{k}\left|f_{G}(u)-f_{G}(v)\right|^{2}\\[8.0pt] &\geq&\left|f_{G}(u)-f_{G}(v)\right|^{2},\end{array}$